04/23/08 – QCE had to fix the same AHU unit again when they cleaned up some of the school panel wiring. Project is nearing completion.

04/11/08 – MAH released the Const Mtg 05 Notes to the project partners.

03/13/08 – (3) full size sets of drawings and a letter stating that no loads were being added and that the feeders and panels are of sufficient size were delivered to IBI today. QCE should be able to obtain a permit with these actions.

03/06/08 – Today’s construction meeting has been rescheduled to 03/17 at 2pm. We also discussed the revised schedule for the project which I updated in a pseudo meeting minute memo in PDF format. TEI will process the IBI letter for no load increases during this job for QCE to give to IBI to get permit squared away. Bert also asked for a revised scheduled update which we have included in the memo.

02/26/08 – QCE called today to tell us that IBI denied our permits for this job. They now instead want TEI to provide a letter saying no loads have been changed or will change during this project and we are doing a direct panel for panel replacement. A shutdown is going to be scheduled in late March for school panelboard replacement(22 or 23).

01/09/08 – Today we decided that the pre const meeting notes were correct and the project completion date for this project is March 1, 2008. QCE will do and update project schedule with delivery details for TEI to approve and give to the owner very soon.

01/07/08 – MAH released the Const Mtg 02 Notes to the project partners.

11/28/07 – QCE provided an update to Mark Donnelly today with the following information: The ID portion of the project is going well. The (4) cottages are complete. Chapel, John Morrow building, Goundskeeping building, & Pole Barn are completed also with the ID. We still have to change locks and install the new locks on certain panels. Administration building is approximately 75% completed, the School is approximately 50% completed. Some work has been done in the Dining building, Rec building, & Gym. We have not been in Security yet. The transformer pad for the school transformer was poured 11/27/07. All Panels have been measured & verified. The order has been placed, but I don’t have delivery dates yet. As soon as I have the delivery dates I will put a schedule together and ask for a meeting to go thru what has to happen for the shutdowns & installation of the panels. TTC setup a const mtg for 12/10 at 2:30pm to follow up with panel order and schedule.

11/12/07 – TTC released the Pre Const Mtg Notes to the project partners.

08/31/07 – TEI still awaiting word from HCFD. This project remains in the Contract Phase.

08/17/07 – TEI awaiting word from HCFD in order to move forward with this project.

08/10/07 – TEI spoke with QCE about their bid. They were comfortable with the bid so TEI wrote the contract and forwarded to the Hamilton County Facilities Department. TEI awaiting word on signed contract so this project can move into the construction phase.

07/31/07 – TEI has obtained a copy of the bid package and is reviewing the information. Dick Kuhn of Queen City put together the bid; he will be back in his office on Wednesday, August 1. NR will contact him in the morning to discuss the bid.

07/27/07 – Today was the bid opening. NR attended for TEI. QCE and Ginter are the low bidders. QCE was at $138,000 for this project. TEI will check the bid to make sure everything is covered.

06/29/07 – No change.

06/22/07 – TEI obtained an ITB #099-07. This project is set to be issued on June 26, 2007. The pre-bid meeting is scheduled for July 3, 2007.

06/15/07 – TEI conducted a site visit on June 13, 2007. The final design strategies have been discussed and this project is moving into its final stages before being issued for bid. TE plans to issue this project for bid by next Friday, June 22 of 2007.

06/08/07 – This project was reviewed in a meeting between TEI and Mark Donnelly. There will be a set of drawings and a cost estimate delivered to Hamilton County Facilities Department this afternoon (June 8, 2007). A meeting will be held on site next week to review final conditions and then will go out for bid.

06/01/07 – NR will have a review set of drawings, specifications and a cost estimate ready for Mark Donnelly to review by the end of next week (June 8, 2007).

05/25/07 – Project is moving along as expected.

05/22/07 – Project review meeting was held today. Various points were discussed as described in the meeting notes. TEI is moving forward with this project while keeping the points discussed in today’s meeting in mind.

05/21/07 – TE has completed all field work. Drawings are 98% complete. A meeting to discuss the status of this project is scheduled for May 22, 2007 with Mark Donnelly.

04/07/07 – TEI has completed the device identification portion of this project and are currently inputting the data into CAD. The panel replacement portion of this project is moving along as expected.

04/20/07 – TEI has been on site working on the device identification portion of the project. The only areas left to complete are north and south wings of the School Building. TEI will be on site today, April 20 2007, to work on the panel replacement portion of this project.

04/13/07 – TEI will be on site beginning next week to work on the electrical device identification portion of this project. Design for panel replacement is moving along as expected.

04/06/07 – Project moving along as expected.

03/28/07 – Kick-off meeting held on site. Buildings to be included in the Device Identification portion of this phase are Administration, School, Chapel, Security, Maintenance, Storage, Dining & Recreation and Gym & Pool. Efforts are aiming for August of 2007 for the shut-down and transformer pad replacement to coincide with the week that school is out of session. Attendees of this meeting include: Mark Donnelly, Bert Watts, Chris Hohmeister, Phil Lambing, Nick Richmond and Ron from Facilities. The next meeting is scheduled for May 2nd at 10:00 on location.

03/23/07 – Design meeting to be held the last week of March 07.

03/20/07 – TE to assist in deciding scope on this project and settle on which buildings are left to do on this site.

03/15/07 – SOSS received and posted to the website. A kickoff meeting will soon be scheduled.

03/14-07 – NR created project journal and filled in back data from previous meetings. Project awaiting signed SOSS to be returned to TE.